OpenStudy (anonymous):

solve the simultaneous equation x+y=2 and x(square)+2y(square)=11

OpenStudy (anonymous):

guys please aswer mee

OpenStudy (anonymous):

from first equation y=2-x put this in the second \[x^2+2(2-x)^2=11\]this is a quadratic for x

OpenStudy (anonymous):

and so wuts the final ans?

OpenStudy (callisto):

@sarah_98 What have you got when you solve the quadratic equation in x?

OpenStudy (anonymous):

i ddnt solve it

OpenStudy (anonymous):

i need someone to solve the whole thing show me the steps and give the final answer

OpenStudy (callisto):

Why don't you try to do it first? Do you know how to solve quadratic equations? By the way, we are supposed to guide you to reach the answer yourself, not giving you all the steps and final answer to your question.

OpenStudy (anonymous):

ya i noe how to factorise

OpenStudy (anonymous):

but im rly bad at it

OpenStudy (callisto):

Have you learnt quadratic formula?

OpenStudy (anonymous):

yes -b+- square root of b square minus 4ac divided by 2a

OpenStudy (callisto):

Right.. Now consider the equation \(x^2 + 2(2-x)^2 = 11\) Expand the second term first, what have you got?

OpenStudy (anonymous):

I solved it and this what ve got 1.72 or 0.387.to get the above answer you wil have to make x the subject of d formula and subtitute x in equation 2' then expand,u wil get a quadratic equation then solve it to get d final result.